A spaceship travelling between the earth and the moon.the distance between the center of the earth and the moon is 3.8x108m.calculate the distance of spaceship from the earth when it experiences zero gravitational force.(mass of the earth=5.97x1024kg;mass of the moon =7.35x1022kg)


1
Expert's answer
2021-01-26T17:49:19-0500

In order to spaceship experiences zero gravitational force, the gravitational forces exerted on the spaceship by the Earth and Moon must be equal:


"F_E=F_M,""\dfrac{GM_EM_{ship}}{r^2}=\dfrac{GM_MM_{ship}}{(R_{EM}-r)^2},""\dfrac{M_E}{r^2}=\dfrac{M_M}{(R_{EM}-r)^2}."

From this equation we can find the distance of spaceship from the Earth when it experiences zero gravitational force:


"r=\dfrac{R_{EM}\sqrt{\dfrac{M_E}{M_M}}}{(1+\sqrt{\dfrac{M_E}{M_M}})},""r=\dfrac{3.8\cdot10^8\ m\cdot\sqrt{\dfrac{5.97\cdot10^{24}\ kg}{7.35\cdot10^{22}\ kg}}}{(1+\sqrt{\dfrac{5.97\cdot10^{24}\ kg}{7.35\cdot10^{22}\ kg}})}=3.4\cdot10^8\ m."